Reports 1-1 of 1 Clear search Modify search
masahide.tamaki - 8:48 Friday 23 December 2022 (23334) Print this report
TF measurement (from MN to DARM and each coil to DARM)

I measured TFs from MN to DARM for each Type-A suspension (DARM / SUSPENSION_MN_DOF_SUMOUT).
Previously, I measured up to 50 Hz (klog22907), but this time it is up to 100 Hz.

In this time, I measured TFs from each coil to DARM to estimate DAC noise (DARM / SUSPENSION_{MN, IM, TM}_COILOUTF_{V1-3, H1-3 for MN&IM, H1~4 for TM}).

There are so many TFs to measure and locking FPMI takes more time than usual (GR was not stable), so I couldn't finish to measure all of them.
The rest are 16 TFs from EY coils to DARM ({MN,IM}_{V1-3, H1-3} & TM_{H1-4}).

I'll post noise calculation after EY measurement log.

All files are in /users/tamaki/low_noise/noise_projection/20221222.

Comments to this report:
masahide.tamaki - 0:31 Saturday 24 December 2022 (23355) Print this report

I measured the remaining 16 TFs (from EY coil out to DARM).
The files are in the same directory as above.

I will post noise caluculation when it is finished (maybe tomorrow).

Search Help